if a^3+b^3+c^3-3abc=0, does it mean that a+b+c=0?
Answer by Edwin McCravy(20086)

if a^3+b^3+c^3-3abc=0, does it mean that a+b+c=0?

No, because if a = b = c = 1 then 

a³ + b³ + c³ - 3abc = 1³ + 1³ + 1³ - 3(1)(1)(1) = 0

yet

a + b + c = 1 + 1 + 1 = 3, not 0.

Edwin